OpenStudy (anonymous):

Solve: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

2x^2+7x+3=0

OpenStudy (anonymous):

i think its {-1/2, -3} but idk

OpenStudy (anonymous):

hmmm..you should get 7 by adding and 3 as product

OpenStudy (nurali):

2x2 + 7x + 3 = 0 2x2 + 6x +x + 3 = 0 2x(x+3)+1(x+3)=0 (2x+1)(x+3)=0 2x+1=0 or x+3=0 x=-1/2 or x=-3

OpenStudy (anonymous):

so {3/2, 1} ?

OpenStudy (akashdeepdeb):

Get 7 by adding and 6 as product okay?

OpenStudy (akashdeepdeb):

You were right allie!! :)

OpenStudy (anonymous):

what about this one Factor: 36x^2 - 25 ?? is it (6x - 5)(6x - 5)

OpenStudy (anonymous):

YUP

OpenStudy (anonymous):

yayyyyy

OpenStudy (anonymous):

oops no.. its 6x+5)(6x-5) @allie22

OpenStudy (anonymous):

oh okay , thank you!

OpenStudy (anonymous):

:)